OpenStudy (anonymous):

could someone assist me, i cant seem to find any reference in my notes. not searching for answers, more of a guide on how to approach the question. produce the following graph for -2 3}}

OpenStudy (dumbcow):

ok for piecewise, you need to get your endpoints f(-2) = 4-(-2)^2 = 0 f(1) = 4 - 1 = 3 (open) f(1) = 3/2 + 3/2 = 3 (closed) f(3) = 3/2(3) + 3/2 = 6 (closed) f(3) = 3+3 = 6 (open) f(5) = 5+3 = 8 plot these points, then connect them with appropriate shape interval 1 --- parabola interval 2,3 ---- line

OpenStudy (dumbcow):

lines are easy for parabola, draw what the entire graph is , then look at shape in interval of [-2,1] insert that shape into piecewise graph

OpenStudy (anonymous):

thank you @dumbcow. i shall let you know of the outcome

OpenStudy (dumbcow):

ok:)